OpenStudy (anonymous):

8. Simplify the following quotient of complex numbers into the form a + bi. 7-8i / -2 + 8i

OpenStudy (anonymous):

answer or method?

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

man i am a different color every time i post. like the horse in the wizard of oz

OpenStudy (anonymous):

oh no

OpenStudy (angela210793):

hahah :D I like silver :D

OpenStudy (anonymous):

if you just want the answer, cheat http://www.wolframalpha.com/input/?i=%287-8i%29+%2F%28+-2+%2B+8i%29

OpenStudy (anonymous):

-39 - 20i ------------ 34

OpenStudy (anonymous):

i dont like these colors at all lol >.<

OpenStudy (angela210793):

i just don't like pink :S

OpenStudy (anonymous):

if you want the method, use \[\frac{a+bi}{c+di}=\frac{a+bi}{c+di}\times \frac{c-di}{c-di}=\frac{(a+bi)(c-di)}{c^2+d^2}\]

OpenStudy (anonymous):

i was pink, orange, red, maybe puce, who knows. now i am nothing. sometimes chair, sometimes squiggle. jeez

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

joe is red

OpenStudy (anonymous):

with a circle

OpenStudy (angela210793):

hahahhaha....then u must stop answering q so u can b a constant color=red

OpenStudy (anonymous):

i feel plain.

OpenStudy (anonymous):

i am going to stop once i get a job

OpenStudy (anonymous):

LOL

OpenStudy (angela210793):

hahah :D

OpenStudy (anonymous):

or a life, whichever comes first. or my new car.

OpenStudy (anonymous):

when school starts i'll only be on like...once a week o.O

OpenStudy (anonymous):

and that will be for help with my work lolol

OpenStudy (anonymous):

yeah but just don't ask me. ode? pde? never heard of it

OpenStudy (anonymous):

we can all learn it then! while im going through my classes.